OpenStudy (anonymous):

w=2csc (1-3x) first derivative of the function

OpenStudy (anonymous):

\[w'=-2\csc \left( 1-3x \right)\cot \left( 1-3x \right)\left( -3 \right)=6\csc \left( 1-3x \right)\cot \left( 1-3x \right)\]

OpenStudy (anonymous):

ow thanks!! how about// tan (1/4 pi - 1/2x )? my answer was 89/2 sec^2 89/2 x